OpenStudy (anonymous):

El Capitan in Yosemite National Park is 3593 ft high. How long would it take a carabiner to fall freely from the top?

OpenStudy (saifoo.khan):

s = 3593 a = 9.8 t = ? u = 0 v = ? Find the "v" first. \[v^2 = u^2 + 2as\]\[v = \sqrt{0^2+2(9.8)(3593)}\]\[v = 265.37ft/s\]

OpenStudy (saifoo.khan):

Now use this to find time taken. using \[v = u + at\]

OpenStudy (unklerhaukus):

time of free fall \[\sqrt{{2h \over g}}\] h = 3593 ft = 1095 m g = 9.81 ms^-2

OpenStudy (unklerhaukus):

about 15 seconds.